OpenStudy (anonymous):

four times a certain number , when increased by 5, is -27

OpenStudy (texaschic101):

4n + 5 = -27 4n = -27 - 5 4n = -32 n = -32/4 n = - 8

OpenStudy (anonymous):

thank you!